Determine the equation of the plane containing the points P(1,3,2) Q=(-2,0,-2) and R=(1,4,3)


1
Expert's answer
2021-12-14T13:04:10-0500

There points

"\begin{aligned}

P=(1,3,2)\\

Q=(-2,0,2) \\

R=(1,4,3)

\end{aligned}"

So, the vector"\overrightarrow {PQ}" and "\overrightarrow{PR}" will be

"\begin{aligned}
\overrightarrow{PQ} =(-2-1) \hat{\imath}+(0-3) \hat{\jmath}+(2-2) \hat{k}=-3 \hat{i}-3 \hat{\jmath} \\
\overrightarrow{PR}=(1-1) \hat{i}+(4-3) \hat{\jmath}+(3-2) \hat{k}=\hat{\jmath}+\hat{k} \\
\end{aligned}"

So, The vector perpendicular to this two vectors is given by

"\overline{P Q} \times \overline{P R}=\left|\begin{array}{ccc}

\hat{i} & \hat{j} & \hat{k} \\

-3 & -3 & 0 \\

0 & 1 & 1

\end{array}\right|"

So. The equation of plane will be

"\begin{gathered}
-3(x-1)+3(y-3)+(-3)(z-2)=0 \\
\Rightarrow x-y+z=2
\end{gathered}"